ROTATIONAL SPECTROSCOPY OF ETHYLAMINE INTO THE THz

\underline{ZBIGNIEW KISIEL}, ADAM KRASNICKI, Institute of Physics, Polish Academy of Sciences, Al. Lotników 32/46, 02-668 Warszawa, Poland; IVAN R. MEDVEDEV, CHRISTOPHER NEESE, SARAH FORTMAN, MANFRED WINNEWISSER, FRANK C. DE LUCIA, Department of Physics, The Ohio State University, Columbus, OH 43210; HOLGER S. P. MÜLLER, I. Physikalisches Institut, Universität zu Köln, Zülpicher Str. 77, 50937 Köln, Germany.

Ethylamine is one of the molecules that exhibits a challenging combination of two low frequency motions: inversion and internal rotation. The separation between these modes is fortunately appreciably greater than in the lighter methylamine, and the cm-wave rotational spectra of anti-ethylamine^{a} and gauche-ethylamine^{b} have been assigned some time ago. Recent astrophysical interest prompted laboratory investigation of anti-ethylamine up to 270 GHz and an associated astronomical search for this species in Sgr B2(N).^{c}

We report extensive new coverage of the rotational spectrum of ethylamine obtained in the form of three broadband segments, each recorded by using a different instrumental technique. The spectrum in the 115-376 GHz region was obtained with the FASSST spectrometer,^{d} at 555-650 GHz by using cascaded harmonic multiplication from a cm-wave synthesizer, and at 867-1081 GHz by also using FASSST, but with frequency tripled output from a 300 GHz-region BWO oscillator. The spectrum is at an advanced stage of analysis, carried out by means of the graphical assignment AABS package,^{e} and results obtained for both anti- and gauche-ethylamine are presented.

aE.Fischer, E.Botskor, J. Mol. Spectrosc., 91, 116-127 (1982).

bE,Fischer, E.Botskor, J. Mol. Spectrosc., 104, 226-247 (1984).

cA.J.Apponi, et al., Astrophys. J., 673, 1240-1248 (2008).

dI.Medvedev et al., J. Mol. Spectrosc., 228, 314-328 (2004).

eZ.Kisiel et al., J. Mol. Spectrosc., 233, 231-243 (2005).